# Effective Malware Detection for Embedded Computing Systems with Limited Exposure

### v1

@@ Abstract
One of the pivotal security threats for the embedded computing systems is malicious software a.k.a malware. With efficiency and efficacy, Machine Learning (ML) has been widely adopted for malware detection in recent times. Despite being efficient, the existing techniques require a tremendous number of benign and malware samples for training and modeling an efficient malware detector. Furthermore, such constraints limit the detection of emerging malware samples due to the lack of sufficient malware samples required for efficient training. To address such concerns, we introduce a code-aware data generation technique that generates multiple mutated samples of the limitedly seen malware by the devices. Loss minimization ensures that the generated samples closely mimic the limitedly seen malware and mitigate the impractical samples. Such developed malware is further incorporated into the training set to formulate the model that can efficiently detect the emerging malware despite having limited exposure. The experimental results demonstrates that the proposed technique achieves an accuracy of 90% in detecting limitedly seen malware, which is approximately 3 $\times$ more than the accuracy attained by state-of-the-art techniques .
@@
Hardware security, malware, machine learning, deep learning, image processing, GANs
@@ I Introduction
  • Introduces a code-aware data generation architecture for increasing the training dataset.
@@ II Related Work
Later [27] introduced a technique for malware detection using image processing where binary applications are converted into grayscale images. The generated images have identical patterns because of the executable file structural distributions. The paper used the K-Nearest Neighbour ML algorithm for the classification of malware images. Other approaches [9, 8] include image visualization and classification using machine learning algorithms such as SVM. However, these approaches don’t address the problem of classifying newer complex malware that is code obfuscated, polymorphed, etc. Neural networks such as ANNs are used extensively to solve the problem [8], as neurons can capture the features of the images accurately than other machine learning algorithms. But, the fully connected layers of artificial neural networks tend to exhaust computational resources. In [28, 29, 30] authors used Convolutional neural networks for malware confinement, as they are popular for their ability to efficiently handle image data through feature extraction by Convolutional 2D layers and using Maxpooling 2D layers to downsample the input parameters, thus, reducing the computational resources. The drawback here is, they need to be trained with large amounts of training data to perform classification efficiently.
@@ VI Conclusion
With the proposed code-aware data generation technique we were able to employ few-shot learning to efficiently classify, limitedly seen malware data. From the results presented, it is evident that the model trained on data, generated through the proposed technique outperforms, the model trained on limited available data, with approximately 9% more accuracy. We also furnished the ASIC implementations of different classifiers trained using the proposed technique. Thus, instead of training the ML models with limited available data, the proposed code-aware data generation technique should be employed.

### v2

# Generative AI-Based Effective Malware Detection for Embedded Computing Systems

@@ Abstract
One of the pivotal security threats for the embedded computing systems is malicious software a.k.a malware. With efficiency and efficacy, Machine Learning (ML) has been widely adopted for malware detection in recent times. Despite being efficient, the existing techniques require a tremendous number of benign and malware samples for training and modeling an efficient malware detector. Furthermore, such constraints limit the detection of emerging malware samples due to the lack of sufficient malware samples required for efficient training. To address such concerns, we introduce a code-aware data generative AI technique that generates multiple mutated samples of the limitedly seen malware by the devices. Loss minimization ensures that the generated samples closely mimic the limitedly seen malware and mitigate the impractical samples. Such developed malware is further incorporated into the training set to formulate the model that can efficiently detect the emerging malware despite having limited exposure. The experimental results demonstrates that the proposed technique achieves an accuracy of 90%, which is approximately 9% higher while training classifiers with only limitedly available training data.
@@
Hardware security, malware, machine learning, deep learning, image processing, Generative AI, GANs
@@ I Introduction
  • Introduces a code-aware generative-AI architecture for increasing the training dataset.
@@ II Related Work
Later [27] introduced a technique for malware detection using image processing where binary applications are converted into grayscale images. The generated images have identical patterns because of the executable file structural distributions. The paper used the K-Nearest Neighbour ML algorithm for the classification of malware images. Other approaches [9, 8] include image visualization and classification using machine learning algorithms such as SVM. However, these approaches don’t address the problem of classifying newer complex malware that is code obfuscated, polymorphed, etc. Neural networks such as ANNs are used extensively to solve the problem [8], as neurons can capture the features of the images accurately than other machine learning algorithms. But, the fully connected layers of artificial neural networks tend to exhaust computational resources. In [28, 29, 30, 31, 32, 33, 23, 22, 34, 35, 36, 37, 38, 39] authors used Convolutional neural networks for malware confinement, as they are popular for their ability to efficiently handle image data through feature extraction by Convolutional 2D layers and using Maxpooling 2D layers to downsample the input parameters, thus, reducing the computational resources. The drawback here is, they need to be trained with large amounts of training data to perform classification efficiently.
@@ VI Conclusion
With the proposed code-aware generative AI technique we were able to employ few-shot learning to efficiently classify, limitedly seen malware data. From the results presented, it is evident that the model trained on data, generated through the proposed technique outperforms, the model trained on limited available data, with approximately 9% more accuracy. We also furnished the ASIC implementations of different classifiers trained using the proposed technique. Thus, instead of training the ML models with limited available data, the proposed code-aware data generation technique should be employed.